OpenStudy (anonymous):

if sin a 15/17 what is cos a

OpenStudy (luigi0210):

\[\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile } \]

OpenStudy (anonymous):

Thank you

OpenStudy (luigi0210):

And as for your question, let's start of by drawing a triangle and labeling it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

This is what I have

OpenStudy (luigi0210):

Do you know what sin is? In terms of sides and such?

OpenStudy (anonymous):

No not in the terms of a triangle

OpenStudy (luigi0210):

\[\sin \theta = \frac{opposite}{hyp}\]

OpenStudy (anonymous):

okay

OpenStudy (luigi0210):